OpenStudy (anonymous):

Bradley and Kelly are out flying kites at a park one afternoon. A model of Bradley and Kelly's kites are shown below on the coordinate plane as kites BRAD and KELLY, respectively: Which statement is correct about the two kites?

OpenStudy (anonymous):

http://gyazo.com/477b395318278036824a90ec23d7bf34

OpenStudy (bibby):

what are the statements?

OpenStudy (anonymous):

a. They are not similar because segment BR to segment KE is 1:2 and segment YL to segment DA is 2:1. b. They are similar because segment BR to segment KY is 1:2 and segment BD to segment YL is 1:2. c. They are not similar because segment YL to segment DA is 1:2 and segment KY to segment BD is 2:1. d. They are similar because segment KY to segment BD is 2:1 and segment YL to segment DA is 2:1.

OpenStudy (bibby):

I think we have to use the distance formula \(d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\) BR=\(\sqrt{(4-3)^2+(6-7)^2}=\sqrt{(1)^2+(-1)^2}=\sqrt2\) KE=\(\sqrt{(11-9)^2+(2-0)^2}=\sqrt{2^2+2^2}=\sqrt{4+4}=\sqrt8\) A is wrong

OpenStudy (bibby):

\(BR=\sqrt{(4-3)^2+(6-7)^2}=\sqrt{(1)^2+(-1)^2}=\sqrt2\) \(KY=\sqrt{(11-9)^2+{2-4}^2}=\sqrt{2^2+({-2})^2}=\sqrt{4+4}=\sqrt8\) B is wrong

OpenStudy (bibby):

\(\sqrt{8}=2\sqrt{2}\)

OpenStudy (bibby):

go with d

OpenStudy (anonymous):

tytytytty. c: ♥

OpenStudy (bibby):

I'll pretend those are hearts

OpenStudy (anonymous):

It was supposed to be only one heart but it made all those. lmao

OpenStudy (bibby):

♥ here's one more heart http://puu.sh/cIJP4.png

OpenStudy (anonymous):

c: